# Strong log-convexity of genus sequences

### v1

@@ 1 Introduction
The main results of this paper are Theorems 3 and 8 that refute the LCGD Conjecture in a rather spectacular (and unexpected) way. Let us state it first in a weaker form.
@@ 2 Multiple log-convexity of genus sequences is viable
Combinatorially, $C$ is a non-separating cycle of the graph $G$ if the subgraph $G-V(C)$ is connected; the cycle $C$ is an induced cycle of $G$ if it is induced as the subgraph of $G$, i.e. it has no chords in $G$. Realizing that facial cycles of 3-connected plane graphs (i.e. graphs of 3-dimensional convex polyhedra) are induced and non-separating, Tutte [44] called such cycles peripheral. Due to this relationship, we say that a 2-cell embedding of a graph in a surface is a polyhedral embedding if every face is bounded by a peripheral cycle.
@@
More generally, we say that a family $\mathcal{C}$ of peripheral cycles in $G$ is peripheral if any two cycles in the family are disjoint and nonadjacent and $G-\cup\{V(C)\mid C\in\mathcal{C}\}$ is connected.
@@
The proof of the lemma can be found in [31] (see also [35, Proposition 4.2.6]). The bound of $23g-$ 2 can actually be improved to $33g-$ 3 if $g\geq 2$ (which is best possible).
@@
Figure 1: (a) The antiprismatic path of length 9 gives the antiprism $\widehat{C}_{9}$ after identifying the left edge $v_{0}^{0}v_{0}^{1}$ with the one on the right. (b) The stacked antiprism $\widehat{C}_{9}^{4}$ of height 4. (c) The genus 0 cylinder $H_{m}^{k,b,d}$ with parameters $m=9$, $k=2$, $b=2$, and $d=3$. The vertices and edges on the left are identified with the vertices and edges on the right side to give a planar graph on the cylinder.
Next, we define the stacked antiprism $\widehat{C}_{m}^{b}$ of height $b$ by stacking $b$ copies of $\widehat{C}_{m}$ on the top of each other (identifying the cycle $v_{1}^{0}v_{2}^{0}\dots v_{m}^{0}$ of the bottom one with the cycle $v_{1}^{1}v_{2}^{1}\dots v_{m}^{1}$ of the copy at the top of it). See Figure 1(b) for an example.
Now, we take the disjoint union of $k+1$ stacked antiprisms $\widehat{C}_{m}^{b}$, where $b$ is a large positive integer (to be specified later), and where $m\geq 6$ is a multiple of 3. For $j=1,\dots,k$, we join the $j$th stacked antiprism with three antiprismatic paths of length $d$ (the value to be specified later) to the $(j-1)$st stacked antiprism. These antiprismatic paths will be referred to as the connectors. The connectors start and end on the stacked antiprisms equally spaced (at distance $\tfrac{1}{3}m-1$ from each other). See Figure 1(c) for an example. The resulting graph $H_{m}^{k,b,d}$ has a natural embedding into the cylinder, and is therefore planar. It is easy to see that $H_{m}^{k,b,d}$ is 4-connected, so its embedding in the plane is polyhedral. There are two faces $F_{1},F_{2}$ of length $m$, one at the top and one at the bottom, and for each $j\in\{1,\dots,k\}$, there are $3$ faces of length $\tfrac{2}{3}m+2d-2$. All other faces are of length $3$.
In order to obtain $G_{g,k}$ for $g\geq 0$, we start with $H_{m}^{k,b,d}$ and fill in each of its faces $F_{1}$ and $F_{2}$ of length $m$ with a dense triangulation such that the following properties hold:
@@
    $G_{g,k}$ contains $g$ isometric non-surface-separating cycles whose pairwise distance and their distance from the subgraph $H_{k}$ is at least $f(g,k)$.
Such triangulations are easy to construct. One can start with any triangulation whose maximum degree is at most 7 and contains a triangle $T_{0}$ whose degrees are all 4. then subdivide each edge and connect the three subdivision vertices in each triangle except $T_{0}$ to subdivide that triangle into four smaller triangles. This can be repeated so long that conditions needed for (a)–(f) are obtained. We omit the details.
Note that for every $g\geq 0$ and $k\geq 1$, $G_{g,k}$ is 4-connected and polyhedrally embedded in the surface of genus $g$. Herewith we will also assume that $m\geq f(g,k)$ and that $b=f(g,k)$. Let $\nu(g,k)$ be the number of vertices of $G_{g,k}$ not counting the vertices on the $3k$ connectors in $H_{m}^{k,b,d}$, i.e.
@@
This completes the description of the graphs $G_{g,k}$. The constructed polyhedral embedding of $G_{g,k}$ has genus $g$ and it will be denoted as $\Pi_{0}$.
@@
Let us now consider 2-cell embeddings of $G_{g,k}$ of genus $h=g+r$, where $1\leq r\leq 2k$. Let us first give a lower bound on $a_{h}(G_{g,k})$.
@@
Let us consider the embedding of $G_{g,k}$ of genus $g+r$ that is obtained from the embedding $\Pi_{0}$ as follows. First, choose $q$ indices $j$ between $1$ and $k$ and for each such $j$ consider the two edges of the connectors on the $j$th stacked antiprism $\widehat{C}_{m}^{b}$. Now change the local rotation at the endpoints of these two edges so that these edges would emanate into the incident big faces. This change merges four triangular faces of $\Pi_{0}$ containing these two edges into two faces of length 4 and merges the three big faces into a single face. Hence the effect of this change is that the genus increases by two. Let $W_{i},W_{i}^{\prime}$ be the two paths on the $i$th connector ($i=1,2,3$), oriented from “bottom” to “top” (see Fig. 2). Then the big new face $F$ contains these paths in the following order: $F=W_{1}^{-}e_{1}W_{1}^{\prime}\cdots W_{2}^{-}e_{2}W_{2}^{\prime}\cdots W_{3}^{-}\cdots W_{3}^{\prime}\cdots$, where $W_{i}^{-}$ denotes the traversal of the path in the opposite direction and $e_{i}$ is the flipped edge (see Fig. 2). Now, each of the $2d$ edges connecting $W_{i}$ and $W_{i}^{\prime}$ can stay where it is or it can be flipped and reembedded into $F$, giving $2^{2d}$ different choices. By doing this for $i=1,2,3$ and for each of the $q$ chosen connectors, we get $2^{6dq}$ different embeddings, all of genus $g+2q$. This proves the claim when $r$ is even.
If $r$ is odd, we select one of the remaining connecting triples and in that case flip the first edge of just one of the connectors. This will further increase the genus by one. In the same way as above we see that we can independently reembed each of the $2d$ edges between $W_{1}$ and $W_{1}^{\prime}$ into the new big face $F$ which has $W_{1}$ and $W_{1}^{\prime}$ on its boundary traversed in the opposite directions. This gives additional $2^{2d}$ possible choices to each of the previously described embeddings of genus $g+r-1$. ∎
@@
We split the vertices of $G=G_{g,k}$ into four sets, $V_{A},V_{B},V_{C},V_{D}$, which are defined as follows. First, $V_{D}$ contains for each of the $3k$ connectors the six vertices in the two triangles containing the first and the last edge of the connector (respectively) and the vertex adjacent to that edge but outside of the connector; the set $V_{C}$ contains the remaining $22d-$ 2 vertices on each of the connectors; the set $V_{B}$ contains all vertices that are incident to any of the nontriangular faces and are not in $V_{C}\cup V_{D}$; lastly, $V_{A}$ contains all remaining vertices. Note that $|V_{D}|=18k$, $|V_{C}|=6k(d-1)$, $|V_{B}|=2k(m-6)$, and $|V_{A}|=\nu(g,k)-2k(m+3)$.
@@
Finally, suppose that $1\leq r<2k$, where $r$ is odd. Then, using Claim 1 for $r-1$ and $r+1$ (and noting that $r-1$ and $r+1$ are even), we see that
@@ 3 Many consecutive log-convex terms
for $1\leq t<k$. This growth is realized by joining $t+1$ large faces of the last connectors where we have $12k+$ 1 connectors of height $c_{2k}d=d$. However, $a_{g+k},a_{g+k+1},\dots,a_{g+2k}$ grow exponentially faster. We claim that
@@
To prove that the right hand side of (5) is an upper bound, we have to prove that using the increase of the genus by $k+t$ as in the lower bound construction above is best possible. To show this we have to argue that merging $k+t+1$ big faces with the highest connectors of height $c_{k+t+1}d$ gives much smaller numbers, and that using the gain where we merge $k+s$ faces of height $c_{k+s}d$ and adding the rest of the genus by merging $t-s+1$ faces with the connectors of height $c_{2k}d$ also gives lower numbers. This comparison reduces to the following two inequalities:
@@ 4 In conclusion
Let us end with a more positive viewpoint. Our counterexamples to the LGCD Conjecture are still unimodal. As far as the author can tell, the unimodal weakening of the LGCD Conjecture could (and actually should) still be true.

### v2

@@ 1 Introduction
The main results of this paper are Theorems 3 and 9 that refute the LCGD Conjecture in a rather spectacular (and unexpected) way. Let us state it first in a weaker form.
@@ 2 Multiple log-convexity of genus sequences is viable
Combinatorially, $C$ is a non-separating cycle of the graph $G$ if the subgraph GV(C)G-V(C) is connected; the cycle $C$ is an induced cycle of $G$ if it is induced as the subgraph of $G$, i.e. it has no chords in $G$. Realizing that facial cycles of 3-connected plane graphs (i.e. graphs of 3-dimensional convex polyhedra) are induced and non-separating, Tutte [44] called such cycles peripheral. Due to this relationship, we say that a 2-cell embedding of a graph in a surface is a polyhedral embedding if every face is bounded by a peripheral cycle.
@@
More generally, we say that a family $\mathcal{C}$ of peripheral cycles in $G$ is peripheral if any two cycles in the family are disjoint and nonadjacent (i.e., no two vertices on distinct cycles are adjacent) and $G-\cup\{V(C)\mid C\in\mathcal{C}\}$ is connected.
@@
The proof of the lemma can be found in [31] (see also [35, Proposition 4.2.6]). The bound of 3g23g-2 can actually be improved to 3g33g-3 if $g\geq 2$ (which is best possible).
@@
Figure 1: (a) The antiprismatic path of length 9 gives the antiprism $\widehat{C}_{9}$ after identifying the left edge $v_{0}^{0}v_{0}^{1}$ with the one on the right. (b) The stacked antiprism $\widehat{C}_{9}^{4}$ of width 9 and height 4. (c) The genus 0 cylinder $H_{m}^{k,b,d}$ with parameters $m=9$, $k=2$, $b=2$, and $d=3$. The vertices and edges on the left are identified with the vertices and edges on the right side to give a planar graph on the cylinder.
Next, we define the stacked antiprism $\widehat{C}_{m}^{b}$ of width $m$ and height $b$ by stacking $b$ copies of $\widehat{C}_{m}$ on the top of each other (identifying the cycle $v_{1}^{0}v_{2}^{0}\dots v_{m}^{0}$ of the bottom one with the cycle $v_{1}^{1}v_{2}^{1}\dots v_{m}^{1}$ of the copy at the top of it). See Figure 1(b) for an example.
Figure 2: (i) Replacing a triangular face by the triangular grid of side length $l$, illustrated here with $l=7$. (ii) Replacing four triangles inside the triangular grid so that a face with all vertices of degree 4 is obtained.
A graph that is 2-cell embedded in a surface $S$ is a near-triangulation if all its faces with the exception of a single face $F$ are triangles. The exceptional face $F$ is called the outer face. It can be of any length, but each edge on $F$ is contained in one of the triangular faces. If $k\geq 1$, we define similarly a $k$-near-triangulation of a surface if there are precisely $k$ faces that are not required to be of length 3.
Lemma 8.
Let $g\geq 0$ and $a,b\geq 1$ be integers. There exists a near-triangulation $T_{g}$ with the following properties:
  • (1)
    The embedding of $T_{g}$ is in the orientable surface of genus $g$.
  • (2)
    The graph $T_{g}$ is 4-connected.
  • (3)
    The maximum vertex degree in $T_{g}$ is at most 7 and any two vertices of degree $7$ are at least distance $a$ apart. The outer face has length $3a$ and each of its vertices has degree $4$.
  • (4)
    $T_{g}$ contains a peripheral family of noncontractible cycles $D_{1},\dots,D_{g}$ that are pairwise at distance at least $2b$, all are of length $3a$, all of their vertices have degree $6$, and each cycle $D_{i}$ passes transversally through each of its vertices, i.e., the two edges of $D_{i}$ incident with a vertex $v$ are opposite in the local rotation around $v$. Moreover, the vertices at distance at most $b$ from $D_{i}$ form a triangulation of the cylinder isomorphic to the stacked antiprism $\widehat{C}_{3a}^{2b}$.
  • (5)
    By cutting the surface along the cycles $D_{1},\dots,D_{g}$ we obtain a surface of genus 0 (with $2g$ boundary components corresponding to the cut).
Proof.
We start with a $5$-connected triangulation of the plane, whose maximum degree is 6 and has at least $2g+1$ faces. Now, we replace each triangle with the triangular grid of side length $5$ (see Figure 2, where the triangular grid of side length $7$ is shown). In $2g+1$ of these triangular grids, we replace the four triangles in the middle with the graph shown in Figure 2. By doing this, we obtain a triangulation of the plane of maximum degree 7 that contains $2g+1$ triangles $F_{0},F_{1},\dots,F_{2g}$, whose vertices are all of degree 4. Now, we once again replace every triangle except $F_{0},\dots,F_{2g}$ by the triangular grid of side length $a$. This gives rise to a $(2g+1)$-near-triangulation, whose outer faces $F_{0}^{\prime},\dots,F_{2g}^{\prime}$ correspond to $F_{0},\dots,F_{2g}$, they are all of length $3a$ and all their vertices are of degree 4.
Finally, for each $i\in\{1,\dots,g\}$, we take a copy of the stacked antiprism $\widehat{C}_{3a}^{2b}$ of width $3a$ and height $2b$ and identify its top and bottom face (each of length $3a$) with $F_{2i-1}^{\prime}$ and $F_{2i}^{\prime}$, respectively. This operation introduces $g$ handles and produces a near-triangulation of the orientable surface of genus $g$ (with outer face $F_{0}^{\prime}$). The cycle of length $3a$ in the middle of the stacked antiprism is then denoted by $D_{i}$.
It is now easy to see that the obtained near-triangulation satisfies all stated conditions. ∎
With this lemma, we are ready to describe the construction for Theorem 3.
@@
In order to construct the desired embedded graph $G_{g,k}$, we first take the disjoint union of $k+1$ stacked antiprisms $\widehat{C}_{m}^{b}$, where $b$ is a large positive integer (to be specified later), and where $m\geq 6$ is a multiple of 3. For $j=1,\dots,k$, we join the $j$th stacked antiprism with three antiprismatic paths of length $d$ (the value to be specified later) to the (j1)(j-1)st stacked antiprism. These antiprismatic paths will be referred to as the connectors. The connectors start and end on the stacked antiprisms equally spaced (at distance $\tfrac{1}{3}m-1$ from each other). See Figure 1(c) for an example. The resulting graph $H_{m}^{k,b,d}$ has a natural embedding into the cylinder, and is therefore planar. It is easy to see that $H_{m}^{k,b,d}$ is 4-connected, so its embedding in the plane is polyhedral. There are two faces $F_{1},F_{2}$ of length $m$, one at the top and one at the bottom, and for each $j\in\{1,\dots,k\}$, there are $3$ faces of length $\tfrac{2}{3}m+2d-2$. All other faces are of length $3$.
Having described the graph $H_{m}^{k,b,d}$ embedded in the plane, we next fill each of its faces $F_{1}$ and $F_{2}$ of length $m$ with a near-triangulation, whose outer facial cycle is identified with the facial cycle of $F_{1}$ or $F_{2}$, such that the following properties hold:
@@
    $G_{g,k}$ contains a peripheral family of $g$ cycles whose pairwise distance and their distance from the subgraph $H_{k}$ is at least $f(g,k)$.
The two near-triangulations to fill-in the two faces $F_{1}$ and $F_{2}$ are easy to obtain by using Lemma 8 with appropriate parameters. We can take the one in $F_{1}$ to have genus $g$ and the one in $F_{2}$ having genus 0.
Note that for every $g\geq 0$ and $k\geq 1$, the resulting embedded graph $G_{g,k}$ is 4-connected and is polyhedrally embedded in the surface of genus $g$. Herewith we will also assume that $m\geq f(g,k)$ and that $b=f(g,k)$. Let $\nu(g,k)$ be the number of vertices of $G_{g,k}$ not counting the vertices on the $3k$ connectors in $H_{m}^{k,b,d}$, i.e.
@@
This completes the description of the graphs $G_{g,k}$. The constructed polyhedral embedding of $G_{g,k}$ has genus $g$ and it will be denoted by $\Pi_{0}$.
@@
Let us now consider 2-cell embeddings of $G_{g,k}$ of genus $h=g+r$, where $1\leq r\leq 2k$. We will first give a lower bound and then an upper bound on $a_{h}(G_{g,k})$. The lower bound in Claim 1 is of the form $2^{2d(r+q)}$, where $q:=\lfloor r/2\rfloor$. The upper bound is of the form $2^{2d(r+q)}p(n,d)$, where $p$ is a polynomial whose degree is bounded in terms of constants $g$ and $k$ and $n$ is the order of the graph. Observe that our construction gives that $n=O(g^{2}k^{2}d)$. For even $r$, the value of $q$ is the same as for $r+1$. We use the lower bound for r1r-1 and $r+1$ when $r$ is even, in which case $\lfloor(r-1)/2\rfloor\cdot\lfloor(r+1)/2\rfloor>2\lfloor r/2\rfloor$, and this yields log-convexity when $d$ is large enough.
Let us first give a lower bound on $a_{h}(G_{g,k})$.
@@
Let us consider the embedding of $G_{g,k}$ of genus $g+r$ that is obtained from the embedding $\Pi_{0}$ as follows. First, choose $q$ indices $j$ between $1$ and $k$ and for each such $j$ consider the two edges of the connectors on the $j$th stacked antiprism $\widehat{C}_{m}^{b}$. Now change the local rotation at the endpoints of these two edges so that these edges would emanate into the incident big faces. This change merges four triangular faces of $\Pi_{0}$ containing these two edges into two faces of length 4 and merges the three big faces into a single face. Hence the effect of this change is that the genus increases by two. Let $W_{i},W_{i}^{\prime}$ be the two paths on the $i$th connector ($i=1,2,3$), oriented from “bottom” to “top” (see Fig. 3). Then the big new face $F$ contains these paths in the following order: $F=W_{1}^{-}e_{1}W_{1}^{\prime}\cdots W_{2}^{-}e_{2}W_{2}^{\prime}\cdots W_{3}^{-}\cdots W_{3}^{\prime}\cdots$, where $W_{i}^{-}$ denotes the traversal of the path in the opposite direction and $e_{i}$ is the flipped edge, i.e., the local rotation at its end is changed as indicated in Fig. 3 so that the long faces merge as indicated above. Now, each of the $2d$ edges connecting $W_{i}$ and $W_{i}^{\prime}$ can stay where it is or it can be flipped and reembedded into $F$, giving $2^{2d}$ different choices. By doing this for $i=1,2,3$ and for each of the $q$ chosen connectors, we get $2^{6dq}$ different embeddings, all of genus $g+2q$. This proves the claim when $r$ is even.
If $r$ is odd, we select one of the remaining connecting triples and in that case flip the first edge of just one of the connectors. This will further increase the genus by one. In the same way as above we see that we can independently reembed each of the $2d$ edges between $W_{1}$ and $W_{1}^{\prime}$ into the new big face $F$ which has $W_{1}$ and $W_{1}^{\prime}$ on its boundary traversed in the opposite directions. This gives additional $2^{2d}$ possible choices to each of the previously described embeddings of genus g+r1g+r-1. ∎
@@
We split the vertices of $G=G_{g,k}$ into four sets, $V_{A},V_{B},V_{C},V_{D}$, which are defined as follows. First, $V_{D}$ contains for each of the $3k$ connectors the six vertices in the two triangles containing the first and the last edge of the connector (respectively) and the vertex adjacent to that edge but outside of the connector; the set $V_{C}$ contains the remaining 2d22d-2 vertices on each of the connectors; the set $V_{B}$ contains all vertices that are incident to any of the nontriangular faces and are not in $V_{C}\cup V_{D}$; lastly, $V_{A}$ contains all remaining vertices. Note that $|V_{D}|=18k$, $|V_{C}|=6k(d-1)$, $|V_{B}|=2k(m-6)$, and $|V_{A}|=\nu(g,k)-2k(m+3)$.
@@
Finally, suppose that $1\leq r<2k$, where $r$ is odd. Then, using Claim 1 for r1r-1 and $r+1$ (and noting that r1r-1 and $r+1$ are even), we see that
@@ 3 Many consecutive log-convex terms
for $1\leq t<k$. This growth is realized by joining $t+1$ large faces of the last connectors where we have $2k+1$ connectors of height $c_{2k}d=d$. However, $a_{g+k},a_{g+k+1},\dots,a_{g+2k}$ grow exponentially faster. We claim that
@@
To prove that the right hand side of (5) is an upper bound, we have to prove that using the increase of the genus by $k+t$ as in the lower bound construction above is best possible. To show this we have to argue that merging $k+t+1$ big faces with the highest connectors of height $c_{k+t+1}d$ gives much smaller numbers, and that using the gain where we merge $k+s$ faces of height $c_{k+s}d$ and adding the rest of the genus by merging ts+1t-s+1 faces with the connectors of height $c_{2k}d$ also gives lower numbers. This comparison reduces to the following two inequalities:
@@ 4 In conclusion
Let us end with a more positive viewpoint. Our counterexamples to the LCGD Conjecture are still unimodal. As far as the author can tell, the unimodal weakening of the LCGD Conjecture could (and actually should) still be true.